Fashionisto Inc. is an upscale clothing store in New York City and London. Each store has two main departments, Men’s Apparel and Women’s Apparel. Marie Phelps, Fashionisto’s CFO, wants to use strategic performance measurement to better understand the company’s financial results. She has decided to use the profit center method to measure performance and has gathered the following information about the two stores and the two departments of the New York City store:

Total net sales $ 4,100,000
Fixed costs
Partly traceable and controllable 280,000
Partly traceable but noncontrollable 250,000
Nontraceable costs 100,000
Total net sales (percent)
London store 40 %
New York store 60
New York—Men’s Apparel 40
New York—Women’s Apparel 60
Cost of goods sold—variable (percent of sales)
London store 52 %
New York—Men’s Apparel 45
New York—Women’s Apparel 60
Variable operating costs (percent of sales)
London store 36 %
New York—Men’s Apparel 22
New York—Women’s Apparel 32
Fixed controllable costs—partly traceable (percent of total)
London store 40 %
New York total 40
Men’s Apparel 45
Women’s Apparel 40
Could not be traced to Men’s or Women’s Apparel 15
Could not be traced to London or New York 20
Fixed noncontrollable costs—partly traceable (percent of total)
London store 50 %
New York total 40
Men’s Apparel 35
Women’s Apparel 15
Could not be traced to Men’s or Women’s Apparel 50
Could not be traced to London or New York 10

Required:

1. Using this information, prepare a contribution income statement for Fashionisto.
Answer #1
Combined Company Not Allowed London Store New York Store Not Allowed Men's Wear Women's Wear
Sales $4,100,000 $0 $1,640,000 $2,460,000 $0 $984,000 $1,476,000
Variable Costs:
     Cost of Goods Sold $2,181,200 $0 $852,800 $1,328,400 $0 $442,800 $885,600
     Operating Costs $1,279,200 $0 $590,400 $688,800 $0 $216,480 $472,320
Total Variable Costs $3,460,400 $0 $1,443,200 $2,017,200 $0 $659,280 $1,357,920
Contribution Margin $639,600 $0 $196,800 $442,800 $0 $324,720 $118,080
Traceable & Controllable Fixed Costs $280,000 $56,000 $112,000 $112,000 $16,800 $50,400 $44,800
Traceable & Noncontrollable Fixed Costs $250,000 $25,000 $125,000 $100,000 $50,000 $35,000 $15,000
Segmental Margin $109,600 ($81,000) ($40,200) $230,800 ($66,800) $239,320 $58,280
Common Fixed Costs $100,000
Net Income (Loss) $9,600
Combined Company Not Allowed London Store New York Store Not Allowed Men's Wear Women's Wear
Sales 4100000 =B2-D2-E2 =B2*40% =B2*60% 0 =E2*40% =E2*60%
Variable Costs:
     Cost of Goods Sold =SUM(C4:E4) 0 =D2*52% =SUM(F4:H4) 0 =G2*45% =H2*60%
     Operating Costs =SUM(C5:E5) 0 =D2*36% =SUM(F5:H5) 0 =G2*22% =H2*32%
Total Variable Costs =SUM(B4:B5) =SUM(C4:C5) =SUM(D4:D5) =SUM(E4:E5) =SUM(F4:F5) =SUM(G4:G5) =SUM(H4:H5)
Contribution Margin =B2-B6 =C2-C6 =D2-D6 =E2-E6 =F2-F6 =G2-G6 =H2-H6
Traceable & Controllable Fixed Costs =SUM(C8:E8) =280000*20% =280000*40% =280000*40% =E8*15% =E8*45% =E8*40%
Traceable & Noncontrollable Fixed Costs =SUM(C9:E9) =250000*10% =250000*50% =250000*40% =E9*50% =E9*35% =E9*15%
Segmental Margin =B7-B8-B9 =C7-C8-C9 =D7-D8-D9 =E7-E8-E9 =F7-F8-F9 =G7-G8-G9 =H7-H8-H9
Common Fixed Costs 100000
Net Income (Loss) =B10-B11